Find value of 'a' that the such points (4,a) and (8,3a) lie on a line with slope m= 1/3
Answer by stanbon(75887)

Find value of 'a' that the such points (4,a) and (8,3a) lie on a line with slope m= 1/3
-----
(3a-a)/(8-4) = 1/3
------
2a/4 = 1/3
----
a/2 = 1/3
a = 2/3
------------
Cheers,
Stan H.
-----------